OpenStudy (anonymous):

Solve the equation for y: 4x - 2y = 6. A. y = 2x + 3 B. y = -2x - 3 C. y = 2x - 3 D. y = -2x + 3

OpenStudy (lgbasallote):

hmm..first thing you might wanna do is isolate y..but remember to keep it positive..since y is negative in its position now..transpose it..so.. 4x - 2y = 6 4x = 6 + 2y 4x - 6 = 2y can you see what to do next now?